# Leveraging differentiable programming in the inverse problem of neutron stars

### v1

@@ I Introduction
[…]
@@
Matter at the most extreme densities, up to $6-8n_{\rm{sat}}$, is only realized inside NSs, the remnants of core-collapse supernovae Lattimer and Prakash (2005). Their macroscopic properties, such as mass, radius, quadrupole moment, and tidal deformability, are determined through the Tolman-Oppenheimer-Volkoff (TOV) equations, which require the EOS for closure. Consequently, measurements of NS properties can then be used to infer the EOS Lindblom (1992), which is referred to as the inverse problem of NSs. Several such measurements are now available. For instance, precise radio timing of binary pulsar systems determines the masses of NSs Hobbs et al. (2006); Demorest et al. (2010). Furthermore, pulse-profile modeling of X-ray emissions from hotspots of millisecond pulsars provides joint estimates on the mass and radius of NSs Watts (2019), and has been achieved with the Neutron star Interior Composition ExploreR (NICERGendreau et al. (2016); Bogdanov et al. (2019); Riley et al. (2023, 2019); Miller et al. (2019a); Riley et al. (2021); Miller et al. (2021). Additionally, the tidal deformability of NSs leaves a measurable imprint on gravitational waves emitted during the inspiral and merger of binary neutron star (BNS) systems. This way, GW170817 Abbott et al. (2017, 2019), the first BNS merger observed by Advanced LIGO Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015), provided the first estimates on NS radii derived from GWs Abbott et al. (2018).
@@
Constraining the EOS with input from nuclear physics and astrophysics is commonly done with Bayesian inference Raithel et al. (2017); Radice et al. (2018); Annala et al. (2018); Raithel et al. (2018); Coughlin et al. (2019); Capano et al. (2020); Miller et al. (2020); Tews et al. (2019); Li et al. (2021); Legred et al. (2021); Huth et al. (2022); Biswas (2022); Raaijmakers et al. (2021); Pang et al. (2021); Biswas (2021); Essick et al. (2021); Pang et al. (2023); Essick et al. (2020a); Brandes et al. (2023); Rose et al. (2023); Brandes and Weise (2024); Rutherford et al. (2024); Biswas and Rosswog (2024); Huang et al. (2024); Raaijmakers et al. (2025); Kashyap et al. (2025); Wu et al. (2025) (see Ref. Koehn et al. (2024a) for a recently compiled overview of EOS constraints). However, current inference methods remain computationally expensive, particularly when directly sampling a set of EOS parameters, as this requires solving the TOV equations on the fly for millions of samples. This computational burden will hinder our ability to extract precise EOS constraints from NS data delivered by future experiments, such as upcoming X-ray telescopes (e.g., eXTP and STROBE-X Zhang et al. (2019); Watts et al. (2019); Ray et al. (2019)) and the next-generation of ground-based GW detectors (i.e., the Einstein Telescope Punturo et al. (2010); Maggiore et al. (2020); ET steering committee et al. (2020, 2024) and Cosmic Explorer Evans et al. (2021)). For example, next-generation GW observatories are projected to detect approximately $\mathcal{O}(10^{5})$ BNS mergers annually Maggiore et al. (2020); Piro et al. (2022). These experiments will measure NS radii potentially with an unprecedented accuracy of $10-100$ meters Chatziioannou (2022); Puecher et al. (2023); Iacovelli et al. (2023); Huxford et al. (2024); Walker et al. (2024). This increase in both the quantity and quality of observations highlights the need for more efficient inference methods for the inverse problem of NSs. Ideally, such methods should be capable of solving the TOV equations on the fly to ensure accurate and robust EOS constraints.
@@
One possible solution, thoroughly explored in past works, is to resort to machine-learning surrogates. For instance, previous works have accelerated the TOV solver by replacing it with machine learning-based approximations using regression methods Richter and Li (2023); Imam et al. (2024a), neural networks Ferreira and Providência (2021); Thete et al. (2023); Tiwari and Pai (2024); Reed et al. (2024); Magnall et al. (2024); Di Clemente et al. (2025); Somasundaram et al. (2024), support vector machines Ferreira and Providência (2021), Gaussian processes and reduced basis methods Reed et al. (2024), and dynamic mode decomposition Lalit et al. (2024) (see Ref. Zhou et al. (2024) for a review). Other works, rather than relying on traditional sampling techniques, use machine learning models to directly recover the EOS from NS data, e.g., with neural networks Fujimoto et al. (2018, 2020); Morawski and Bejger (2020); Traversi and Char (2020); Krastev (2022); Fujimoto et al. (2021); Soma et al. (2022, 2023a); Farrell et al. (2023); Han et al. (2023); Ferreira et al. (2022); Carvalho et al. (2023); Soma et al. (2023b); Krastev (2023); Wu and Wen (2024); Chatterjee et al. (2024), normalizing flows Morawski and Bejger (2022); Brandes et al. (2024); Hu et al. (2024), transformers Gonçalves et al. (2023), principal component analysis Patra et al. (2023), and symbolic regression Imam et al. (2024b); Patra et al. (2025). However, training these surrogate models can take on the order of hours McGinn et al. (2024) to weeks Hu et al. (2024) (depending on the model complexity and hardware used for training) and requires additional modeling efforts. Moreover, we are often interested in comparing inferences with different parameterizations or prior assumptions. Since emulators must be retrained whenever these change, their computational speed-up comes at the cost of reduced flexibility in inference.
@@ II.1 Equation of state
Several approaches exist to parametrize the EOS without resorting to a specific physical description of supra-nuclear matter. These include piecewise polytropes Read et al. (2009); Steiner et al. (2010); Hebeler et al. (2013); Raithel et al. (2016); O’Boyle et al. (2020), the spectral representation Lindblom (2010); Lindblom and Indik (2012); Lindblom (2018); Fasano et al. (2019); Lindblom (2022), a speed-of-sound parametrization Tews et al. (2018b); Greif et al. (2019), and non-parametric models based on Gaussian processes Landry and Essick (2019); Essick et al. (2020b); Landry et al. (2020) or neural networks Han et al. (2021). In this work, we adopt the parametrization of Ref. Koehn et al. (2024a).
At the lowest densities, up to $0.5n_{\rm{sat}}$, the crust EOS is fixed to the model from Ref. Douchin and Haensel (2001). At densities just above the crust-core transition, we assume that matter consists solely of nucleonic degrees of freedom in $\beta$-equilibrium and use a metamodel (MM) parametrization of the EOS Margueron et al. (2018a, b); Somasundaram et al. (2021). This parametrizes the energy per nucleon
$\begin{split}e(n,\delta)=e_{\rm{sat}}(n)+e_{\rm{sym}}(n)\delta^{2}+\mathcal{O}(\delta^{4})\,,\end{split}$ (1)
@@
$\begin{split}e_{\rm{sat}}(n)=E_{\rm{sat}}+\frac{1}{2}K_{\rm{sat}}x^{2}+\frac{1}{3!}Q_{\rm{sat}}x^{3}+\frac{1}{4!}Z_{\rm{sat}}x^{4}+\dots\,,\end{split}$ (2)
@@
The expansion coefficients in Eqs. (2) and (3) are referred to as the nuclear empirical parameters. The MM parametrization allows us to incorporate constraints from nuclear-physics experiments or compare them with similar constraints obtained from astrophysical observations. For instance, the slope of the symmetry energy $L_{\rm{sym}}$ is correlated with both the neutron-skin thickness as well as the radii of NSs Typel and Brown (2001); Lattimer and Prakash (2016); Russotto et al. (2023).
However, this parametrization, describing nucleonic degrees of freedom, is inadequate to model the higher-density regime probed by NSs, as new degrees of freedom might become relevant Lattimer and Prakash (2006); Oertel et al. (2016); Kumar et al. (2023), possibly with a phase transition Baym et al. (2018); Pang et al. (2020); Bogdanov et al. (2022); Mondal et al. (2023); Essick et al. (2023); Prakash et al. (2024); Huang and Sourav (2025). Therefore, we use a phenomenological and model-agnostic approach at higher densities. Specifically, we use a speed-of-sound extension scheme (CSE) parametrization above a breakdown density $n_{\rm{break}}$, which is varied freely in our parametrization within $1-2\ n_{\rm{sat}}$ Tews et al. (2018b), unless stated otherwise. At densities above $n_{\rm{break}}$, we parametrize the speed-of-sound $c_{s}$ using a non-uniform grid of points in the $(n,c_{s}^{2})$-plane and construct the full speed-of-sound profile with linear interpolation. We extend this CSE to $25n_{\rm{sat}}$, the end-point of our EOSs, for which $c_{s}^{2}$ is also sampled. The speed-of-sound curve from the CSE parametrization is then matched to that of the MM parametrization with cubic splines. The EOS can then be derived from the speed-of-sound curve, see Ref. Somasundaram et al. (2023).
@@ II.2 Neutron star observables
The macroscopic properties of non-spinning, isolated, spherically symmetric NSs are determined through the TOV equations Tolman (1939); Oppenheimer and Volkoff (1939), the general-relativistic equations determining hydrostatic equilibrium [^1] We refer readers to Ref. Glendenning (2010) for a didactic derivation of the TOV equations.:
@@
where $\varepsilon$ is the energy density, $r$ is the radial coordinate from the center of the NS and $m(r)$ is the mass contained in a sphere of radius $r$. Here, we employ units where $G=c=1$. The system of equations is closed by the EOS, which links pressure and energy density. Given a central pressure $p_{c}$ of the NS, the TOV equations can be integrated outwards, with boundary conditions ${m(r=0)=0}$ and ${p(r=0)=p_{c}}$, until we reach the boundary of the star where ${p(r=R)=0}$. This gives the star’s radius $R$ and gravitational mass ${M\equiv m(R)}$. Repeating this procedure for various $p_{c}$ provides the mass-radius relation of NSs for that EOS.
When an NS is placed in a static, external quadrupolar tidal field $\mathcal{E}_{ij}$, it develops a quadrupole moment $Q_{ij}$ Hinderer et al. (2010); Damour and Nagar (2009); Damour et al. (2012). This induced deformation is described by the $\ell=2$ eigenfunctions of the oscillations and, to linear order, takes the form $Q_{ij}=-\lambda\mathcal{E}_{ij}$, where $\lambda$ is related to the second tidal Love number $k_{2}$ through $k_{2}=\frac{3}{2}\lambda R^{-5}$. We refer readers to Refs. Hinderer et al. (2010); Damour and Nagar (2009); Damour et al. (2012) for further details. These tidal deformations occur, for instance, in BNS systems, where the gravitational field of each NS induces the tidal effects deforming its companion. This leaves an imprint on the phase of the GWs emitted during inspiral. In particular, the GW emission depends, at leading order, on each star’s dimensionless tidal deformability parameter $\Lambda$, which is defined as
@@ II.3 Differentiable programming
Differentiable programming is a programming paradigm that enables the end-to-end differentiation of computer programs Baydin et al. (2018); Sapienza et al. (2024); Blondel and Roulet (2024). The derivative is computed with automatic differentiation, which exploits the fact that all numerical calculations are composed of elementary operations. The derivatives of these operations are accumulated during runtime to efficiently evaluate the gradient of a program.
@@
In this work, we use jax Frostig et al. (2018); Kidger (2021), which provides the aforementioned automatic differentiation capabilities in Python. Moreover, jax can compile Python code and execute it on hardware accelerators, such as a GPU. To exploit these features, we have developed a Python package implementing jax-based EOS code and a TOV solver, called jester [^2] https://github.com/nuclear-multimessenger-astronomy/jester. With jester, we can efficiently evaluate the derivative of any numerical function that depends on the EOS parameters, even if it involves solving the TOV equations to obtain the macroscopic properties of the NS as an intermediate step. In the following sections, we introduce how we leverage automatic differentiation in gradient-based samplers and a novel gradient-based inversion scheme for NSs.

### v2

@@ I Introduction
The nuclear EOS remains uncertain and an open problem in physics Lattimer:2012nd; Ozel:2016oaf; Burgio:2021vgk; Chatziioannou:2024tjq. Nuclear experiments only constrain its lower density behavior, i.e., around the nuclear saturation density, $\rho_{\rm{sat}}\equiv 2.7\times 10^{14}\ \rm{g}\ \rm{cm}^{-3}$} (or equivalently, a baryon number density of $n_{\rm{sat}}\equiv$0.16\text{\,}{\mathrm{fm}}^{-3}$\$}). For instance, experiments measuring the neutron-skin thickness of atomic nuclei constrain the symmetry energy at nuclear saturation Birkhan:2016qkr; Thiel:2019tkm; PREX:2021umo; Reed:2021nqk; CREX:2022kgg; Giacalone:2023cet, while heavy-ion collisions uniquely probe isospin-symmetric matter above the nuclear saturation density Danielewicz:2002pu; Li:2008gp; Huth:2021bsp; Sorensen:2023zkk. In addition, nuclear theory calculations provide constraints on isospin-asymmetric matter up to two times the nuclear saturation density Tews:2018kmu; Drischler:2020hwi; Armstrong:2025tza.
@@
Matter at the most extreme densities, up to $6-8n_{\rm{sat}}$, is only realized inside NSs, the remnants of core-collapse supernovae Lattimer:2004sa. Their macroscopic properties, such as mass, radius, quadrupole moment, and tidal deformability, are determined through the Tolman-Oppenheimer-Volkoff (TOV) equations, which require the EOS for closure. Consequently, measurements of NS properties can then be used to infer the EOS Lindblom1992-zb, which is referred to as the inverse problem of NSs. Several such measurements are now available. For instance, precise radio timing of binary pulsar systems determines the masses of NSs Hobbs:2006cd; Demorest:2010bx. Furthermore, pulse-profile modeling of X-ray emissions from hotspots of millisecond pulsars provides joint estimates on the mass and radius of NSs Watts:2019lbs, and has been achieved with the Neutron star Interior Composition ExploreR (NICERgendreau2016neutron; Bogdanov:2019ixe; riley2023xpsi; Riley:2019yda; Miller:2019cac; Riley:2021pdl; Miller:2021qha. Additionally, the tidal deformability of NSs leaves a measurable imprint on gravitational waves emitted during the inspiral and merger of binary neutron star (BNS) systems. This way, GW170817 LIGOScientific:2017vwq; LIGOScientific:2018hze, the first BNS merger observed by Advanced LIGO LIGOScientific:2014pky and Advanced Virgo VIRGO:2014yos, provided the first estimates on NS radii derived from GWs LIGOScientific:2018cki.
@@
Constraining the EOS with input from nuclear physics and astrophysics is commonly done with Bayesian inference Raithel:2017ity; Radice:2017lry; Annala:2017llu; Raithel:2018ncd; Coughlin:2018fis; Capano:2019eae; Miller:2019nzo; Tews:2019cap; Li:2020bre; Legred:2021hdx; Huth:2021bsp; Biswas:2021pvm; Raaijmakers:2021uju; Pang:2021jta; Biswas:2021yge; Essick:2021ezp; Pang:2022rzc; Essick:2020flb; Brandes:2022nxa; Rose:2023uui; Brandes:2024wpq; Rutherford:2024srk; Biswas:2024hja; Huang:2024rfg; Raaijmakers2025; Kashyap:2025cpd; Wu:2025edc (see Ref. Koehn:2024set for a recently compiled overview of EOS constraints). However, current inference methods remain computationally expensive, particularly when directly sampling a set of EOS parameters, as this requires solving the TOV equations on the fly for millions of samples. This computational burden will hinder our ability to extract precise EOS constraints from NS data delivered by future instruments, such as upcoming X-ray telescopes (e.g., eXTP and STROBE-X eXTP:2018anb; Watts:2018iom; STROBE-XScienceWorkingGroup:2019cyd) and the next-generation of ground-based GW detectors (i.e., the Einstein Telescope Punturo:2010zza; ET:2019dnz; ETScienceCase2020; ETDesignUpdate2024 and Cosmic Explorer Evans:2021gyd). For example, next-generation GW observatories are projected to detect approximately $\mathcal{O}(10^{5})$ BNS mergers annually ET:2019dnz; Piro:2021oaa. These detectors will measure NS radii potentially with an unprecedented accuracy of 1010010-100 meters Chatziioannou:2021tdi; Puecher:2023twf; Iacovelli:2023nbv; Huxford:2023qne; Walker:2024loo. This increase in both the quantity and quality of observations highlights the need for more efficient inference methods for the inverse problem of NSs. Ideally, such methods should be capable of solving the TOV equations on the fly to ensure accurate and robust EOS constraints.
@@
One possible solution, thoroughly explored in past works, is to resort to machine-learning surrogates. For instance, previous works have accelerated the TOV solver by replacing it with machine learning-based approximations using regression methods Richter:2023zec; Imam:2023ngm, neural networks Ferreira:2019bny; Thete:2022drz; Tiwari:2024jui; Reed:2024urq; Magnall:2024ffd; DiClemente:2025pbl; Somasundaram:2024ykk, support vector machines Ferreira:2019bny, Gaussian processes and reduced basis methods Reed:2024urq, and dynamic mode decomposition Lalit:2024vmu (see Ref. Zhou:2023pti for a review). Other works, rather than relying on traditional sampling techniques, use machine learning models to directly recover the EOS from NS data, e.g., with neural networks Fujimoto:2017cdo; Fujimoto:2019hxv; Morawski:2020izm; Traversi:2020dho; Krastev:2021reh; Fujimoto:2021zas; Soma:2022qnv; Soma:2022vbb; Farrell:2022lfd; Han:2022sxt; Ferreira:2022nwh; Carvalho:2023ele; Soma:2023fyi; Krastev:2023fnh; Wu:2023npy; Chatterjee:2023ecc; Ventagli:2024xsh, normalizing flows Morawski:2022aud; Brandes:2024vhw; Hu:2024oen, transformers Goncalves:2022smd, principal component analysis Patra:2023jbz, and symbolic regression Imam:2024xqg; Patra:2025sak; Patra:2025xtd; Bejger:2025hbq. However, training these surrogate models can take on the order of hours McGinn:2024nkd to weeks Hu:2024oen (depending on the model complexity and hardware used for training) and requires additional modeling efforts. Moreover, we are often interested in comparing inferences with different parameterizations or prior assumptions. Since emulators must be retrained whenever these change, their computational speed-up comes at the cost of reduced flexibility in inference.
@@ II.1 Equation of state
Several approaches exist to parametrize the EOS without resorting to a specific physical description of supra-nuclear matter. These include piecewise polytropes Read:2008iy; Steiner:2010fz; Hebeler:2013nza; Raithel:2016bux; OBoyle:2020qvf, the spectral representation Lindblom:2010bb; Lindblom:2012zi; Lindblom:2018rfr; Fasano:2019zwm; Lindblom:2022mkr, a speed-of-sound parametrization Tews:2018iwm; Greif:2018njt, and non-parametric models based on Gaussian processes Landry:2018prl; Essick:2019ldf; Landry:2020vaw or neural networks Han:2021kjx; Li:2025obt. In this work, we adopt the parametrization of Ref. Koehn:2024set.
At the lowest densities, up to $0.5n_{\rm{sat}}$, the crust EOS is fixed to the model from Ref. Douchin:2001sv. At densities just above the crust-core transition, we assume that matter consists solely of nucleonic degrees of freedom in $\beta$-equilibrium and use a metamodel (MM) parametrization of the EOS Margueron:2017eqc; Margueron:2017lup; Somasundaram:2020chb. This parametrizes the energy per nucleon
[…]
@@
Throughout this work, we fix $E_{\rm{sat}}=-$16\text{\,}\mathrm{MeV}$\$ and $n_{\rm{sat}}=0.16\ \rm{fm}^{-3}$. The expansion of the symmetry energy is given by
@@
The expansion coefficients in Eqs. (2) and (3) are referred to as the nuclear empirical parameters. The MM parametrization allows us to incorporate constraints from nuclear-physics experiments or compare them with similar constraints obtained from astrophysical observations. For instance, the slope of the symmetry energy $L_{\rm{sym}}$ is correlated with both the neutron-skin thickness as well as the radii of NSs Typel:2001lcw; Lattimer:2015nhk; Russotto:2023ari.
However, this parametrization, describing nucleonic degrees of freedom, is inadequate to model the higher-density regime probed by NSs, as new degrees of freedom might become relevant Lattimer:2006qiu; Oertel:2016xsn; Kumar:2023lhv, possibly with a phase transition Baym:2017whm; Pang:2020ilf; Bogdanov:2022faf; Mondal:2023gbf; Essick:2023fso; Prakash:2023afe; Huang:2025vfl. Therefore, we use a phenomenological and model-agnostic approach at higher densities. Specifically, we use a speed-of-sound extension scheme (CSE) parametrization above a breakdown density $n_{\rm{break}}$, which is varied freely in our parametrization within $1-2\ n_{\rm{sat}}$ Tews:2018iwm, unless stated otherwise. At densities above $n_{\rm{break}}$, we parametrize the speed-of-sound $c_{s}$ using a non-uniform grid of points in the $(n,c_{s}^{2})$-plane and construct the full speed-of-sound profile with linear interpolation. We extend this CSE to $25n_{\rm{sat}}$, the end-point of our EOSs, for which $c_{s}^{2}$ is also sampled. The speed-of-sound curve from the CSE parametrization is then matched to that of the MM parametrization with cubic splines. The EOS can then be derived from the speed-of-sound curve, see Ref. Somasundaram:2021clp.
@@ II.2 Neutron star observables
The macroscopic properties of non-spinning, isolated, spherically symmetric NSs are determined through the TOV equations Tolman:1939jz; Oppenheimer:1939ne, the general-relativistic equations determining hydrostatic equilibrium [^1] We refer readers to Ref. glendenning2010special for a didactic derivation of the TOV equations.:
@@
where $\varepsilon$ is the energy density, $r$ is the radial coordinate from the center of the NS and $m(r)$ is the mass contained in a sphere of radius $r$. Here, we employ units where $G=c=1$. The system of equations is closed by the EOS, which links pressure and energy density. Given a central pressure $p_{c}$ of the NS, the TOV equations can be integrated outwards, with boundary conditions m(r=0)=0{m(r=0)=0} and $p(r=0)=p_{c}$}, until we reach the boundary of the star where p(r=R)=0{p(r=R)=0}. This gives the star’s radius $R$ and gravitational mass $M\equiv m(R)$}. Repeating this procedure for various $p_{c}$ provides the mass-radius relation of NSs for that EOS.
When an NS is placed in a static, external quadrupolar tidal field $\mathcal{E}_{ij}$, it develops a quadrupole moment $Q_{ij}$ Hinderer:2009ca; Damour:2009vw; Damour:2012yf. This induced deformation is described by the $\ell=2$ eigenfunctions of the oscillations and, to linear order, takes the form $Q_{ij}=-\lambda\mathcal{E}_{ij}$, where $\lambda$ is related to the second tidal Love number $k_{2}$ through $k_{2}=\frac{3}{2}\lambda R^{-5}$. We refer readers to Refs. Hinderer:2009ca; Damour:2009vw; Damour:2012yf for further details. These tidal deformations occur, for instance, in BNS systems, where the gravitational field of each NS induces the tidal effects deforming its companion. This leaves an imprint on the phase of the GWs emitted during inspiral. In particular, the GW emission depends, at leading order, on each star’s dimensionless tidal deformability parameter $\Lambda$, which is defined as
@@ II.3 Differentiable programming
Differentiable programming is a programming paradigm that enables the end-to-end differentiation of computer programs Baydin:2015tfa; sapienza2024differentiableprogrammingdifferentialequations; blondel2024elementsdifferentiableprogramming. The derivative is computed with automatic differentiation, which exploits the fact that all numerical calculations are composed of elementary operations. The derivatives of these operations are accumulated during runtime to efficiently evaluate the gradient of a program.
@@
In this work, we use jax frostig2018compiling; kidger2021on, which provides the aforementioned automatic differentiation capabilities in Python. Moreover, jax can compile Python code and execute it on hardware accelerators, such as a GPU. To exploit these features, we have developed a Python package implementing jax-based EOS code and a TOV solver, called jester [^2] https://github.com/nuclear-multimessenger-astronomy/jester. With jester, we can efficiently evaluate the derivative of any numerical function that depends on the EOS parameters, even if it involves solving the TOV equations to obtain the macroscopic properties of the NS as an intermediate step.
While the TOV boundary condition (pressure reaching zero) appears non-differentiable, our implementation circumvents this challenge through two key techniques. First, we parameterize the TOV equations using enthalpy $h$ as the integration variable (similar to, e.g., Ref. Lindblom1992-zb), which transforms the stopping condition from p(r)=0{p(r)=0} to the smooth endpoint h=0{h=0}. Second, we pre-allocate arrays for a maximum number of integration steps and employ bounded while loops to populate the solution arrays, after which jax operations mask and truncate the results to extract only the physically meaningful portion. A similar strategy is employed to truncate the NS solutions at the TOV mass. Since these operations rely entirely on jax primitives, the entire code supports automatic differentiation.
In the following sections, we introduce how we leverage automatic differentiation in gradient-based samplers and a novel gradient-based inversion scheme for NSs.
@@ II.4 Bayesian inference
It has previously been shown that flowMC generates more effective samples than an MCMC algorithm that only uses a gradient-based sampler Gabrie:2021tlu. In this work, we use the Metropolis-adjusted Langevin algorithm (MALA) for the gradient-based sampler grenander1994representations. We have not tested our setup with the NUTS sampler, a specific implementation of Hamiltonian Monte Carlo (HMC) Betancourt:2017ebh; Hoffman:2011ukg. However, MALA is computationally cheaper to run (see, e.g., Ref. Neural_Langevin_Dynamical_Sampling for runtime comparisons on some numerical experiments) and gives satisfactory results in terms of effective sample size.